OpenStudy (anonymous):

which graph represents the solution set of x^(2)-2>0?

OpenStudy (vishweshshrimali5):

\[\large{x^2-2>0}\] \[\large{\implies x^2 > 2}\] Now, remember this: \[\large{\text{if}~~ y^2 = a^2}\] \[\large{\text{then}~~ |y| = a}\] Thus, \[\large{x^2 > 2}\] \[\large{\implies|x| > \sqrt{2}}\] \[\large{\implies x > \sqrt{2}~~\text{or}~~ x < -\sqrt{2}}\]  Shaded region represent the inequality.